Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In A configuration (08 Jun); short meeting following the holiday. Martin away this week but will push ahead characterizing data rate behavior (currently still at 15 MB/s). Correlator hardware (64 BlB) all available. EA24 out of array (lightning+). ALMA Early science statistics. Transition to preparing for receiving proposals to submitting them!
  • Key dates:
    • 08 Jun - 12 Sep: A configuration (move from 31 May - 08 Jun).
    • 01 Aug - Proposal Deadline (C, CnB, B)
  • Staff Travel:
    • RP: 10 Jul - 23 Jul (UK, SA)
    • RP: 09 Aug - 21 Aug (URSI)
    • FO: 12 Aug - 12 Sep
    • RP: 10 Sep - 16 Sep (Brazil)
  • Visitors:
    • Robert Harris (AC982): 24 Apr - 16 Jul (High frequency w/ Claire)
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA21 back to array 11 Jul.
    • Receivers: EA12 X band; EA21 spate of receivers by the end of the month (28 Jun); ACU work scoped out in time and resources.
    • Calibration (Perley): Gain compression/expansion still being looked at. (at the site today). Antennas with new RF switches; tests run on this morning.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): Stable weekend; ongoing work with Brent on simulations to understand noise characteristics/V stokes behavior.
  • Computing (Robnett):

Software status and updates (Butler)

  • Links:
  • SSS: Work on new OSRO/RSRO capabilities. OPT testing page ready: https://staff.nrao.edu/evla/sss/opt/testingTips.1.10.00.html (change in database (Oracle to postgress); Correlator: flipping between versions to test data rate issues; operations using stable CBE (can still drop a scan). MCAF update (in support of holography).
  • Data Rate Discussion: Currently promising 25 MB/s; we're capable of approximately 15 MB/s. Work to improve warnings and notification about the limitations.
  • CASA Update (new functionality available/issues/enhancements):Notes: new Autoflagging capabilities.
  • AIPS Update (new functionality available/issues/enhancements): Notes: Eric out today; more next week.
  • SDM Update: Notes: Ephemeris table status.

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Configuration move week (work around these activities:
  • Mon - Holiday
  • Tue
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Identified possible issues in the noise related to the 3-level phase rotator.
  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags). .
  • Correlator: [Clark/Carlson] Delay rocking across bands. Ongoing discussion of implementation; pending Bryan's return
  • Array: [Gustaaf] Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).
  • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week/soon

  • CM/CBE: Data rate testing and push; work on new missing scans.
  • Holography mode testing.
  • V Stokes follow-up.
  • Subarrays: cleanup.
  • 3-bit sampler testing; automatic gain setting.
  • Stage 2 mixer testing to relax sub 128 MHz BW grids
  • HPC large dataset (500 GB)
  • Executor fshift algorithm (new; increments <100 Hz)
  • Phased array/astrometry mode testing.
  • Correlator: Rupen: S-band shadow testing
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Baseband edge/filter rolloff.
  • Sub-band dependent fshifts.
  • Array: *; Individual issue sleuthing (focus on phase jumps).

Commissioning Groups

  • Big deadline is end of this month for incorporation into links for Call for Proposals due 30 June.

    • 4-band (Owen):
      • Notes:
    • Adv: Low band (Owen):
      • Notes:
    • Adv: Astrometry (Mioduszewski):
      • Notes: Begin testing of UVW precision issues (Brunthaler). Report next week.
  • Calibrator Models
    • Notes: A configuration data being collected.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes:

RSRO/ECSO Reports

RSRO A configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011BRSROStatus

Individual reports (updates)

Misc:

Demo Science Status

AOB

Misc Actions

  • Old/Low priority
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • [Rupen]: Development of WIDAR guidelines document

-- JosephMcMullin - 2011-07-05
Topic revision: r1 - 2011-07-05, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback